# IMPACT: Behavioral Intention-aware Multimodal Trajectory Prediction with Adaptive Context Trimming

### v2

@@ 4.1 Experimental Setup
Our experiments are conducted on one of the most challenging prediction datasets, Waymo Open Motion Dataset (WOMD)[39]. This large-scale dataset comprises 486,995 training clips, 44,097 validation clips, and 44920 testing clips. Each clip contains 10 timesteps of historical agent states, 1 current timestep, and 80 future timesteps at a sampling frequency of 10 Hz, along with HD map information. We evaluate our method on both core WOMD tasks: the marginal motion and interactive motion prediction tasks. Following the standard evaluation protocol, we adopt metrics including Soft mAP, mAP, minADE, minFDE, Miss Rate, and Overlap Rate, and Soft mAP is the main metric.
@@
We employ AdamW optimizer [22] for training, conducting experiments on a cluster of 8 NVIDIA A800 GPUs with a total batch size of 80. The learning rate is initialized as $1\times 10^{-4}$ and begins step decay starting at epoch 22, halving every two epochs. The model undergoes 30 epochs. More details can be referred to Appendix section 4.
@@ 4.2 Leaderboard Performance
Figure 5 illustrates the detailed qualitative results for both marginal and joint predictions using our proposed method. For additional qualitative results, please refer to section 7 in Appendix.
RMP-YOLO(Ensemble)[30] 0.4737 0.4531 0.5564 1.1188 0.1084 ${0.1259}^{*}$
ModeSeq(Ensemble)[49] 0.4737 0.4665 0.5680 1.1766 0.1204 0.1275
BeTop[21] 0.4698 0.4587 0.5716 1.1668 0.1183 0.1272
MGTR[10] 0.4599 0.4505 0.5918 1.2135 0.1298 0.1275
EDA[19] 0.4596 0.4487 0.5718 1.1702 0.1169 0.1266
MTR++[28] 0.4410 0.4329 0.5906 1.1939 0.1298 0.1281
MTR[27] 0.4403 0.4249 0.5964 1.2039 0.1312 0.1274
HPTR[42] 0.3968 0.3904 0.5565 1.1393 0.1434 0.1366
HDGT[15] 0.3709 0.3577 0.5933 1.2055 0.1511 0.1557
DenseTNT[13] - 0.3281 1.0387 1.5514 0.1573 0.1779
SceneTransformer[24] - 0.2788 0.6117 1.2116 0.1564 0.1473
Ours (Ensemble) 0.4801 ${0.4598}^{*}$ 0.5563 1.1295 0.1087 0.1258
Ours (Single) 0.4721 0.4609 ${0.5641}^{*}$ ${1.1540}^{*}$ ${0.1143}^{*}$ 0.1255
BeTop-ens [21] 0.2573 0.2511 0.9779 2.2805 0.4376 $0.1688^{*}$
BeTop [21] $0.2466^{*}$ $0.2412^{*}$ 0.9744 2.2744 0.4355 0.1696
MTR++ [28] 0.2368 0.2326 $\underline{0.8975}$ $\underline{1.9509}$ 0.4143 0.1665
MTR [27] 0.2078 0.2037 $0.9181^{*}$ 2.0633 0.4411 0.1717
MotionDiffuser [16] 0.2047 0.1952 0.8642 1.9482 $\underline{0.4300}$ 0.2004
GameFormer [14] 0.1982 0.1923 0.9721 2.2146 0.4933 0.2022
BIFF [50] 0.1383 0.1229 0.9388 $2.0443^{*}$ 0.4620 0.2172
DenseTNT [13] - 0.1647 1.1417 2.4904 0.5350 0.2309
M2I [32] - 0.1239 1.3506 2.8325 0.5538 0.2757
SceneTransformer [24] - 0.1192 0.9774 2.1892 0.4942 0.2067
Ours (Single) 0.2718 0.2659 0.9738 2.2734 $0.4316^{*}$ 0.1684
4.3 Ablation Study
Different behavior modeling and map interaction mechanisms.
Table 3 presents an ablation study evaluating various agent-pruning and map-pruning strategies on the validation set. The first row (All + Dynamic) queries all agents for the decoder while employing dynamic map selection as in MTR [27], serving as a baseline. Introducing agent selection via Braid Theory [21] (Row 2) enhances Soft mAP to 0.4602. Replacing Braid Theory with our two-class filtering—where Yielding, Overtaking, and Nearby behaviors are treated as interactive while Ignore is considered non-interactive (Row 3)—further refines minADE to 0.5693 while maintaining a comparable Miss Rate. Notably, adopting our four-class YOIN approach (Row 4) elevates Soft mAP to 0.4728, demonstrating its ability to better capture nuanced interactions. Lastly, incorporating vectorized occupancy into YOIN (Row 5) further improves performance (Soft mAP = 0.4758), highlighting the benefit of explicit map filtering in isolating relevant polylines and refining trajectory predictions. These results validate the effectiveness of our integrated behavior modeling and map selection framework.
Agent Map Metrics
All w/Braid Theory[21] w/Inter. w/YOIN w/Dynamic[27] w/Vect. Occ. Soft mAP mAP minADE Miss Rate Inference Time (ms/scenario) Params (M)
0.4582 0.4414 0.5718 0.1203 14.67 69.469
0.4602 0.4507 0.5745 0.1187 23.92 48.938
0.4663 0.4549 0.5693 0.1187 23.92 48.938
0.4728 0.4613 0.5680 0.1163 23.92 48.941
0.4758 0.4646 0.5652 0.1173 18.27 48.944
@@ 4.3 Ablation Study
Table 1 shows that our method achieves state-of-the-art (SOTA) performance on the Waymo 2024 motion prediction benchmark, outperforming all existing LiDAR-free approaches in the primary metric, Soft mAP, as well as in minADE and Overlap Rate. In the remaining metrics (mAP, minFDE, and Miss Rate), our method ranks second. Here, ’Ensemble’ denotes the model-ensemble technique described in [27] for performance boosting, while ’Single’ indicates evaluation with a single model.
@@
As presented in Table 2, even without any model-ensemble techniques [27], our single model achieves the best performance on the Waymo joint prediction leaderboard, surpassing the previous SOTA method [21] by 10.2% in both Soft mAP and mAP. Furthermore, our model attains the second-lowest Overlap Rate and the third-lowest Miss Rate. These substantial improvements underscore the effectiveness of the IMPACT framework.
@@ 4.5 Vectorized Occupancy Prediction Performance
Table 5 presents precision, recall, and F1-score for both classes, along with their data ratios in the validation set. The occupied class constitutes only 3.96% of samples, highlighting a strong imbalance. Despite this, the model achieves a high overall accuracy of 0.989. For the occupied class, precision (0.933) and recall (0.775) yield an F1-score of 0.847, indicating rare false positives—crucial for pruning irrelevant map polylines in trajectory prediction. The dominant unoccupied class achieves near-perfect precision (0.991), recall (0.998), and F1-score (0.994). These results demonstrate that our vectorized occupancy prediction effectively identifies critical map segments while minimizing false classification. More qualitative results (like Figure 4) are in Appendix section 7.
@@
As shown in Figure 6, selecting 24 agents and 192 polylines achieves the optimal balance between interaction diversity and trajectory accuracy. On average, the model input contains 43.85 agents and 749.52 map polylines, meaning this selection reduces scene complexity by 45.2% and 74%, respectively, while preserving essential contextual cues. Beyond this threshold, redundant elements introduce noise, degrading generalization and increasing prediction uncertainty. This highlights the importance of selective context pruning to enhance trajectory forecasting performance.
@@
We compare various agent-pruning and map-pruning strategies in Table 3 using the same encoder architecture. The first row (All + w/Dynamic) corresponds to a baseline that queries all agents and dynamically collect polylines, resulting in a model with 69.469 M parameters and a reported inference time of 14.67 ms under a simplified setting. By contrast, our final approach (the last row in Table 3) combines YOIN-based agent pruning with vectorized occupancy-based pruning. This design lowers the parameter count to 48.944 M while reducing the average inference time of BeTOP[21] from 23.92 to 18.27 ms per scenario. The key to these efficiency gains lies in pruning irrelevant scene information—both agents and polylines—based on their semantic contributions, which in turn eliminates redundant computations.
Table 6: Performance gains from integrating proposed modules with existing approaches(Argoverse1&2). bFDE6 is main metric.

### v3

@@ 4.1 Experimental Setup
Our experiments are conducted on one of the most challenging prediction datasets, Waymo Open Motion Dataset (WOMD)[waymo_motion_prediction]. This large-scale dataset comprises 486,995 training clips, 44,097 validation clips, and 44920 testing clips. Each clip contains 10 timesteps of historical agent states, 1 current timestep, and 80 future timesteps at a sampling frequency of 10 Hz, along with HD map information. We evaluate our method on both core WOMD tasks: the marginal motion and interactive motion prediction tasks. Following the standard evaluation protocol, we adopt metrics including Soft mAP, mAP, minADE, minFDE, Miss Rate, and Overlap Rate, and Soft mAP is the main metric.
@@
We employ AdamW optimizer [loshchilov2019decoupledweightdecayregularization] for training, conducting experiments on a cluster of 8 NVIDIA A800 GPUs with a total batch size of 80. The learning rate is initialized as $1\times 10^{-4}$ and begins step decay starting at epoch 22, halving every two epochs. The model undergoes 30 epochs. More details can be referred to Appendix.
@@ 4.2 Leaderboard Performance
Figure 4 illustrates the detailed qualitative results for both marginal and joint predictions using our proposed method. LABEL:fig:figure-sampling-space shows our qualitative marginal and joint prediction results. For additional qualitative results, please refer to Appendix.
@@
Table 3 shows that our method achieves state-of-the-art (SOTA) performance on the Waymo 2024 motion prediction benchmark, outperforming all existing LiDAR-free approaches in the primary metric, Soft mAP, as well as in minADE and Overlap Rate. In the remaining metrics (mAP, minFDE, and Miss Rate), our method ranks second. Here, ’Ensemble’ denotes the model-ensemble technique described in [shi2023MTR] for performance boosting, while ’Single’ indicates evaluation with a single model.
@@
As presented in Table 4, even without any model-ensemble techniques [shi2023MTR], our single model achieves the best performance on the Waymo joint prediction leaderboard, surpassing the previous SOTA method [liu2024reasoningmultiagentbehavioraltopology] by 10.2% in both Soft mAP and mAP. Furthermore, our model attains the second-lowest Overlap Rate and the third-lowest Miss Rate. These substantial improvements underscore the effectiveness of the IMPACT framework.
4.3 Behavioral Intention Prediction Performance
@@ 4.4 Vectorized Occupancy Prediction Performance
Table 2 presents precision, recall, and F1-score for both classes, along with their data ratios in the validation set. The occupied class constitutes only 3.96% of samples, highlighting a strong imbalance. Despite this, the model achieves a high overall accuracy of 0.989. For the occupied class, precision (0.933) and recall (0.775) yield an F1-score of 0.847, indicating rare false positives—crucial for pruning irrelevant map polylines in trajectory prediction. The dominant unoccupied class achieves near-perfect precision (0.991), recall (0.998), and F1-score (0.994). These results demonstrate that our vectorized occupancy prediction effectively identifies critical map segments while minimizing false classification. More qualitative results are in Appendix.
RMP-YOLO(Ensemble)[sun2024rmpyolorobustmotionpredictor] 0.4737 0.4531 0.5564 1.1188 0.1084 ${0.1259}^{*}$
ModeSeq(Ensemble)[zhou2024modeseq] 0.4737 0.4665 0.5680 1.1766 0.1204 0.1275
BeTop[liu2024reasoningmultiagentbehavioraltopology] 0.4698 0.4587 0.5716 1.1668 0.1183 0.1272
MGTR[gan2024mgtrmultigranulartransformermotion] 0.4599 0.4505 0.5918 1.2135 0.1298 0.1275
EDA[lin2024eda] 0.4596 0.4487 0.5718 1.1702 0.1169 0.1266
MTR++[shi2024mtr++] 0.4410 0.4329 0.5906 1.1939 0.1298 0.1281
MTR[shi2023MTR] 0.4403 0.4249 0.5964 1.2039 0.1312 0.1274
HDGT[JiaHDGT] 0.3709 0.3577 0.5933 1.2055 0.1511 0.1557
DenseTNT[gu2021densetnt] - 0.3281 1.0387 1.5514 0.1573 0.1779
SceneTransformer[ngiam2022scenetransformer] - 0.2788 0.6117 1.2116 0.1564 0.1473
Ours (Ensemble) 0.4801 ${0.4598}^{*}$ 0.5563 1.1295 0.1087 0.1258
Ours (Single) 0.4721 0.4609 ${0.5641}^{*}$ ${1.1540}^{*}$ ${0.1143}^{*}$ 0.1255
BeTop-ens [liu2024reasoningmultiagentbehavioraltopology] 0.2573 0.2511 0.9779 2.2805 0.4376 $0.1688^{*}$
BeTop [liu2024reasoningmultiagentbehavioraltopology] $0.2466^{*}$ $0.2412^{*}$ 0.9744 2.2744 0.4355 0.1696
MTR++ [shi2024mtr++] 0.2368 0.2326 $\underline{0.8975}$ $\underline{1.9509}$ 0.4143 0.1665
MTR [shi2023MTR] 0.2078 0.2037 $0.9181^{*}$ 2.0633 0.4411 0.1717
MotionDiffuser [jiang2023motiondiffuser] 0.2047 0.1952 0.8642 1.9482 $\underline{0.4300}$ 0.2004
GameFormer [huang2023gameformergametheoreticmodelinglearning] 0.1982 0.1923 0.9721 2.2146 0.4933 0.2022
BIFF [zhu2023biffbilevelfuturefusion] 0.1383 0.1229 0.9388 $2.0443^{*}$ 0.4620 0.2172
DenseTNT [gu2021densetnt] - 0.1647 1.1417 2.4904 0.5350 0.2309
M2I [sun2022m2ifactoredmarginaltrajectory] - 0.1239 1.3506 2.8325 0.5538 0.2757
SceneTransformer [ngiam2022scenetransformer] - 0.1192 0.9774 2.1892 0.4942 0.2067
Ours (Single) 0.2718 0.2659 0.9738 2.2734 $0.4316^{*}$ 0.1684
4.5 Ablation Study
Different behavior modeling and map interaction mechanisms.
Table 5 presents an ablation study evaluating various agent-pruning and map-pruning strategies on the validation set. The first row (All + Dynamic) queries all agents for the decoder while employing dynamic map selection as in MTR [shi2023MTR], serving as a baseline. Introducing agent selection via Braid Theory [liu2024reasoningmultiagentbehavioraltopology] (Row 2) enhances Soft mAP to 0.4602. Replacing Braid Theory with our two-class filtering—where Yielding, Overtaking, and Nearby behaviors are treated as interactive while Ignore is considered non-interactive (Row 3)—further refines minADE to 0.5693 while maintaining a comparable Miss Rate. Notably, adopting our four-class YOIN approach (Row 4) elevates Soft mAP to 0.4728, demonstrating its ability to better capture nuanced interactions. Lastly, incorporating vectorized occupancy into YOIN (Row 5) further improves performance (Soft mAP = 0.4758), highlighting the benefit of explicit map filtering in isolating relevant polylines and refining trajectory predictions. These results validate the effectiveness of our integrated behavior modeling and map selection framework.
Agent Map Metrics
All w/Braid Theory[liu2024reasoningmultiagentbehavioraltopology] w/Inter. w/YOIN w/Dynamic[shi2023MTR] w/Vect. Occ. Soft mAP mAP minADE Miss Rate Inference Time (ms/scenario) Params (M)
0.4582 0.4414 0.5718 0.1203 14.67 69.469
0.4602 0.4507 0.5745 0.1187 23.92 48.938
0.4663 0.4549 0.5693 0.1187 23.92 48.938
0.4728 0.4613 0.5680 0.1163 23.92 48.941
0.4758 0.4646 0.5652 0.1173 18.27 48.944
@@ 4.5 Ablation Study
On average, the model input contains 43.85 agents and 749.52 map polylines, meaning this selection reduces scene complexity by 45.2% and 74%, respectively, while preserving essential contextual cues. Beyond this threshold, redundant elements introduce noise, degrading generalization and increasing prediction uncertainty. This highlights the importance of selective context pruning to enhance trajectory forecasting performance.
@@
We compare various agent-pruning and map-pruning strategies in Table 5 using the same encoder architecture. The first row (All + w/Dynamic) corresponds to a baseline that queries all agents and dynamically collect polylines, resulting in a model with 69.469 M parameters and a reported inference time of 14.67 ms under a simplified setting. By contrast, our final approach (the last row in Table 5) combines YOIN-based agent pruning with vectorized occupancy-based pruning. This design lowers the parameter count to 48.944 M while reducing the average inference time of BeTOP[liu2024reasoningmultiagentbehavioraltopology] from 23.92 to 18.27 ms per scenario. The key to these efficiency gains lies in pruning irrelevant scene information—both agents and polylines—based on their semantic contributions, which in turn eliminates redundant computations. Notably, the approach retains or improves accuracy metrics such as Soft mAP, mAP, and minADE, underscoring that targeted context selection effectively balances computational efficiency with predictive performance.
Table 6: Performance gains from integrating proposed modules with existing approaches (Argoverse 1 & 2). bFDE6 is the main metric.